# Complexity of Linear Subsequences of $k$-Automatic Sequences

### v1

@@ 4.1 The Thue-Morse Sequence
Next we apply Theorem 10 to the Thue-Morse sequence $(t(i))_{i\geq 0}$ and $n=1$.
@@
Unlike the case of Theorem 17, where the construction in Theorem 10 gives the minimal automaton, in the case of Theorem 20, the construction in Theorem 10 does not necessarily give the minimal automaton, so the upper bound in Theorem 20 is not necessarily tight. The exact number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$ for small values of $c$ is given as sequence A382296 of the OEIS [15]. We state the following as an open problem.
@@
What is the exact number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$ in terms of $c$?
@@ 5 Construction by Büchi Arithmetic
Taking into account the minimization steps and the time spent on creating $M_{n}$ and $M_{c}$ based on the proof of theorems 23 and 24, the time spent on creating $M_{n,c}$ is $O(n\log(c)\log(nc)+(n+c)\log(n+c))$. So for large enough values of $n$ and $c$, the time spent on creating $M_{n,c}$ is $O((n+c)\log(c)\log(nc))$. ∎

### v2

@@ 4.1 The Thue-Morse Sequence
We now turn to shifts of the Thue-Morse sequence. We apply Theorem 10 to the Thue-Morse sequence $(t(i))_{i\geq 0}$ and $n=1$.
@@
Unlike the case of Theorem 17, where the construction in Theorem 10 gives the minimal automaton, in the case of Theorem 20, the construction in Theorem 10 does not necessarily give the minimal automaton, so the upper bound in Theorem 20 is not tight. The next theorem gives a lower bound on the state complexity of $(t(i+c))_{i\geq 0}$.
Theorem 21.
The minimal msd-first DFAO computing $(t(i+c))_{i\geq 0}$ has more than $c^{0.694}$ states for infinitely many $c$.
Proof.
The idea is to consider the case $c=2^{n}$ for $n\geq 1$. In this case we show, using the Myhill-Nerode theorem, that the minimal DFAO for $(t(i+c))_{i\geq 0}$ has at least $F_{n+1}$ states. We let $S_{n}$ be the set of all binary words of length $n$ that start with 0 and have no two consecutive $1$’s. As is well-known, this set has cardinality $F_{n+1}$. We now show the members of this set are pairwise inequivalent under Myhill-Nerode.
Let $x,y\in S_{n}$ with $x\not=y$. We need to find a word $z$ such that $t([xz]_{2}+2^{n})\not=t([yz]_{2}+2^{n})$. If $t([x]_{2}+2^{n})\not=t([y]_{2}+2^{n})$, we can choose $z=\lambda$, the empty string.
Otherwise, assume $t([x]_{2}+2^{n})=t([y]_{2}+2^{n})$. Since $x,y$ are distinct, there must be some position $p$ (numbered from left to right, starting with index $1$) at which $x[p]\not=y[p]$. Without loss of generality, assume $x[p]=0$ and $y[p]=1$. Choose $z=0^{p}$. Let $x^{\prime}$ denote the binary representation of $xz]_{2}+2^{n}$ and $y^{\prime}$ denote the binary representation of $yz]_{2}+2^{n}$. Letting $|u|_{1}$ denote the number of $1$’s in a word $u$, we clearly have $|x^{\prime}|_{1}=|x|_{1}+1$, since a new $1$ arises from the addition of $2^{n}$. However, $|y^{\prime}|_{1}=|y|_{1}$, since adding $2^{n}$ causes the $1$ at position $p$ to become 0 and causes the 0 at position p1p-1 to become $1$, while no other bits change (since $y$ has no two consecutive $1$’s). It follows that $t([xz]_{2}+2^{n})\not=t([yz]_{2}+2^{n})$, as desired. Hence all the elements of $S_{n}$ are pairwise inequivalent, and so the minimal DFAO for $(t(i+c))_{i\geq 0}$ has $F_{n+1}$ states.
Finally, from the estimate $F_{n+1}\geq{1\over 2}\alpha^{n}$ (which is easily proved by induction), where $\alpha=(1+\sqrt{5})/2\doteq 1.61803$ is the golden ratio, we see that if $c=2^{n}$ then the minimal DFAO has at least
$F_{n+1}\geq{1\over 2}\alpha^{n}={1\over 2}c^{{\log\alpha}\over{\log 2}}\geq c^{0.694}$
states, for $n$ large enough. (Note that $(\log\alpha)/(\log 2)\doteq 0.69424$.) ∎
The exact number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$ for small values of $c$ is given as sequence A382296 of the OEIS [15]. We state the following as an open problem.
@@
What is a good formula for the exact number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$, as a function of $c$?
@@ 5 Construction by Büchi Arithmetic
Taking into account the minimization steps and the time spent on creating $M_{n}$ and $M_{c}$ based on the proof of theorems 24 and 25, the time spent on creating $M_{n,c}$ is $O(n\log(c)\log(nc)+(n+c)\log(n+c))$. So for large enough values of $n$ and $c$, the time spent on creating $M_{n,c}$ is $O((n+c)\log(c)\log(nc))$. ∎